# Ontology-driven Prompt Tuning for LLM-based Task and Motion Planning

### v1

@@ Abstract
Performing complex manipulation tasks in dynamic environments requires efficient Task and Motion Planning (TAMP) approaches, which combine high-level symbolic plan with low-level motion planning. Advances in Large Language Models (LLMs), such as GPT-4, are transforming task planning by offering natural language as an intuitive and flexible way to describe tasks, generate symbolic plans, and reason. However, the effectiveness of LLM-based TAMP approaches is limited due to static and template-based prompting, which struggles in adapting to dynamic environments and complex task contexts. To address these limitations, this work proposes a novel ontology-driven prompt-tuning framework that employs knowledge-based reasoning to refine and expand user prompts with task contextual reasoning and knowledge-based environment state descriptions. Integrating domain-specific knowledge into the prompt ensures semantically accurate and context-aware task plans. The proposed framework demonstrates its effectiveness by resolving semantic errors in symbolic plan generation, such as maintaining logical temporal goal ordering in scenarios involving hierarchical object placement. The proposed framework is validated through both simulation and real-world scenarios, demonstrating significant improvements over the baseline approach in terms of adaptability to dynamic environments, and the generation of semantically correct task plans. Videos and code can be found here:  https://muhayyuddin.github.io/llm-tamp/.
@@ I Introduction
Task and motion planning (TAMP) is an essential component in sequential manipulation to perform tasks in complex environments. TAMP involves two stages: high-level symbolic plan generation, which develops a long-term abstract action sequence, and low-level motion planning, which determines the trajectories with geometric constraints required to execute the computed symbolic plan. The classical solution to TAMP problems consists in describing the task in a specific planning representation, such as the Planning Domain Description Language (PDDL) [1], and using a task planner, such as GraphPlan [2] or FastForward [3], to compute a symbolic plan, whose execution requires the call to a motion planner for all actions involving movements. A custom-designed communication layer is usually used to enable interaction between the task and motion planning modules.
Describing the task specification using PDDL for TAMP has proved very successful. However, PDDL has some drawbacks; it requires manual coding of all real-world descriptions and task constraints, which is impractical in a dynamic environment, and in addition, its syntax may be difficult for new users to understand [4]. On the other hand, compared to PDDL, natural language offers a more intuitive and user-friendly interface for task descriptions [5]. In this line, recent advances in pre-trained large language models (LLMs), such as GPT4 [6], have shown good performance in various natural language tasks, and this has accelerated research on the use of LLMs in task planning [7] and TAMP [8] [9].
Figure 1: Task execution using the ontology-driven LLM-TAMP approach. The prompt is processed through the Prompt Tuning module, which uses knowledge-based reasoning to elaborate the user prompt. The state of the environment is textualized using ontological knowledge by the State Descriptor. This detailed elaboration enables the LLM to generate a semantically correct symbolic plan.
LLM-based methods for TAMP typically use pre-trained LLM models, such as GPT4, that receive task descriptions as input [9]. The LLM module is responsible for computing symbolic plans and reasoning for failures, showing significant capabilities when in-context prompting is provided [10]. However, a weak prompt could drastically reduce planning and reasoning capabilities, and hence several recent studies proposed different ways of prompt engineering for better task elaboration. For example, a template-based prompt is proposed in [9], where fixed templates with variables are used to describe spatial and geometric relations, or the proposal in [8], where a template is used to describe the scene and another to textualize the state of the environment. And to better exploit natural language, an interactive task planning using LLM was proposed in [11], where the user input was integrated with stored task guidelines to generate the prompt.
@@
To address the above-stated challenges, this study proposes an ontology-driven prompt-tuning approach that utilizes knowledge-based reasoning to refine and expand the user’s prompt within the task context. Building on the baseline LLM3 TAMP framework [9], we extend it with our novel prompt-tuning method. Some other approaches, such as [13], have also proposed to integrate ontological knowledge with LLM-based task planning, but their primary focus is on objects properties to define optimal grasping strategies (e.g., glass $\rightarrow$ pick gently, wood $\rightarrow$ pick stably). On the contrary, we employ knowledge-based reasoning for prompt tuning to refine the action sequence for correct temporal goal ordering, integrating it with the user’s prompt input for the LLM. This ensures that the generated plan is not only correct, but also semantically accurate. For example, consider the scenario depicted in Fig. 1, where the user instructs to "put the banana, apple, and bowl in the plate". In most cases, the symbolic baseline planner generates a sequence like "pick banana and apple, and place in the plate" followed by "pick bowl and place in the plate". Although it is syntactically correct, it is semantically flawed, as it could result in placing the bowl on top of the apple and banana. Our proposed prompt-tuning approach addresses this by incorporating domain-specific knowledge. It integrates a guiding note into the prompt, specifying that "put bowl before food items because crockery has priority over food items. Put banana and apple after crockery because food items have less priority." This extended prompt will help the LLM generate a semantically correct and consistent plan.
The core contribution of this work is a novel ontology-driven prompt-tuning framework for task and motion planning that helps LLM to compute semantically accurate plans. Building upon this core framework, the following contributions are highlighted:
  • Automated prompt-tuning with task contextual reasoning: This approach extracts task-related common sense and procedural knowledge from the ontology to guide the large language model (LLM) to generate a semantically accurate and logically ordered plan.
@@ II-A Task and Motion Planning
TAMP uses a structured approach where task planning and motion planning are handled separately but in a coordinated manner. Task planning typically uses symbolic AI techniques, such as hierarchical planning, to generate a sequence of discrete actions based on predefined rules or logic. Motion planning uses geometric algorithms, such as sampling-based planners, to compute feasible paths [14] [15]. A declarative language, such as PDDL, is used for TAMP. However, its use is limited due to the requirement of a large number of parameters to define the problem, actions, and effects [16]. Moreover, as the action space expands, finding geometrically feasible symbolic action sequences becomes computationally challenging without effective heuristics [17]. Recent studies have explored data-driven heuristics to enhance TAMP efficiency [18] [19]. Other approaches use knowledge-based reasoning to reduce the action space, using domain-specific ontological knowledge, and then apply heuristics for efficient symbolic plan generation [20] [21] [22]. The approach proposed here is a general manipulation framework to work across various domains, with the specific application ontology providing the specialized knowledge needed to correctly solve the task.
Figure 2: The ontology-driven LLM-TAMP framework enhances prompt elaboration for generating semantically accurate symbolic plans. It begins by processing the user input to extract actions and objects through semantic tagging. The Contextual Inference Engine uses SPARQL queries to retrieve object types and priorities from the ontology, ensuring the correct action sequence based on predefined rules. The Perception Module, with YOLO-based object detection and FoundationPose for object pose estimation, provides real-time spatial data. This information is textualized using ontological knowledge by the Environmental State Descriptor and fed into the Prompt Generator. The final prompt is then fed into the LLM Task Planner, which produces a structured task plan. Finally, the Motion Planner ensures the robot executes the task with feasible, collision-free movements.
II-B Manipulation planning with LLMs
@@
Large Language Models (LLMs) are increasingly being explored to enhance TAMP by using their ability to understand and generate human-like instructions, descriptions, and planning sequences, i.e., LLMs can be used to translate natural language task descriptions into formal task representations without the need of manually representing the problem domain [23, 24]. Due to their extensive language understanding and contextual reasoning capabilities, LLMs are also used to generate high-level task plans and their translation into feasible motion sequences with minimal additional training [25, 26]. For instance, LLM-GROP [8] utilizes a pre-trained LLM to define symbolic goals for object placement and rearrangement, integrating these with traditional TAMP methods, and AutoTAMP [4] employs a pre-trained LLM to convert user inputs into formal language, which is then used in a TAMP algorithm.
@@
In all the approaches mentioned above, effective prompting techniques are crucial to optimize LLM performance [12]. In addition, [27] identified key challenges in prompting large language models (LLMs) for robotic manipulation, including the need for detailed and accurate problem descriptions, as well as managing biases that shape response structures. This study tackles the issue of effective prompting for task and motion planning by utilizing ontological knowledge to enhance task elaboration for LLM.
@@ III Problem Formulation
This section formally defines the task and motion planning problem and how we model the problem for ontology-driven LLM-TAMP.
@@ III-A Task and Motion Planning
  • $\mathcal{A}$ is the set of primitive actions, each one requiring the call to the motion planner for its execution. A primitive action $a\in\mathcal{A}$ is defined by a set of objects $o\in\mathcal{O}$ and a set of continuous parameters $\gamma$, e.g., Place(cup, [x, y, z, $\theta$]). The parameters $\gamma$ serve as goals for the motion planner, and $a$ is considered feasible if the motion planner finds a solution, i.e., a collision-free trajectory $\tau$. We denote such a feasible action as $a(\tau)$.
@@
The objective in Task and Motion Planning is to derive a sequence of feasible actions ${a_{0}(\tau_{0}),a_{1}(\tau_{1}),\dots,a_{T}(\tau_{T})}$, such that $g(s_{T+1})=1$ and $s_{t+1}=\mathcal{T}(s_{t},a_{t}(\tau_{t}))$, for each $t=0,1,\dots,T$.
@@ III-B Problem Modeling
The ontology-driven LLM-TAMP framework is defined as a tuple $\langle\mathcal{P},\mathcal{K},\xi,\mathcal{X},\mathcal{L}\rangle$, where each component plays a crucial role in structuring a feasible action plan:
@@
  • $\mathcal{K}$ denotes the ontological knowledge base, which provides domain-specific information and task-related rules.
@@
    $\mathcal{L}:\mathcal{P}\times\epsilon\times\mathcal{X}\rightarrow a$ (2)
This model receives the prompt $\mathcal{P}$, enhanced task knowledge $\epsilon$, and TAMP parameters $\mathcal{X}$, and it outputs a set of feasible actions $a\subset\mathcal{A}$. The generated action sequence is expressed as $a=\{a_{0}(\tau_{0}),a_{1}(\tau_{1}),\dots,a_{T}(\tau_{T})\}$, where each action $a_{t}$ with trajectory $\tau_{t}$ enables the robot to transition from the initial environment state to the desired goal state.
@@ IV Proposed approach
This section explains how domain-specific ontological knowledge is used to generate suggestions for the LLM, ensuring semantically correct symbolic plan generation with proper temporal goal ordering. The architecture of our prompt-tuning framework is illustrated in Fig. 2. It is composed of four core modules: the Prompt Construction Layer, Ontological Prompt Tuning Module, Perception Module, and LLM-TAMP Module.
The Prompt Construction Layer comprises three submodules: 1) the User Input which describes the task to be performed, 2) the Prompt Template which contains the static sections of the prompt, provides general task descriptions and structure the overall prompt [^1] Complete template can be found here: https://github.com/Muhayyuddin/llm-tamp/blob/llm-site/assets/prompt/txt, and 3) the Environmental State Descriptor that translates the environmental state into a text format, describing the spatial layout and types of objects in the environment. The subsections below will explain the framework’s module in detail.
IV-A Application ontology: classes, properties and rules
@@
We developed a simple kitchen domain ontology to provide semantic reasoning for performing tasks in a kitchen environment. This domain-specific ontology enhances reasoning about task constraints, priority rules and object relations, and for this, three main classes have been defined: Task, ActionPriority and Object. The Object class is further subdivided into six subclasses describing different kitchen objects types: FoodItems, BoxedFood, Crockery, Utensil, KitchenItems, and Container. The hierarchy of classes, object properties, data properties, and some example individuals are shown in Fig. 3.
@@
We define a set of rules [^2] Complete rule set used in this study is available here: https://github.com/Muhayyuddin/llm-tamp/blob/llm-site/assets/smart_kitchen.rdf that specify the correct order of action sequences when manipulating objects in the kitchen environment. These rules are implemented as instances of ActionPriority, where each rule associates a task (such as put or clean) with an object type (such as FoodItem or Crockery) and assigns a priority level. Through these rules, the reasoning mechanism infers a semantically correct action sequence for task execution. For example, Eq. (IV-A) and (IV-A) show the description logic representation for two rules, named Rule 1 and Rule 2, that state the priorities for the put task: Rule 1 states that the crockery has higher priority and should be placed first, and Rule 2 states that food items have lower priority and should be placed after the crockery.

### v2

# Onto-LLM-TAMP: Knowledge-oriented Task and Motion Planning using Large Language Models

@@ Abstract
Performing complex manipulation tasks in dynamic environments requires efficient Task and Motion Planning (TAMP) approaches that combine high-level symbolic plans with low-level motion control. Advances in Large Language Models (LLMs), such as GPT-4, are transforming task planning by offering natural language as an intuitive and flexible way to describe tasks, generate symbolic plans, and reason. However, the effectiveness of LLM-based TAMP approaches is limited due to static and template-based prompting, which limits adaptability to dynamic environments and complex task contexts. To address these limitations, this work proposes a novel Onto-LLM-TAMP framework that employs knowledge-based reasoning to refine and expand user prompts with task-contextual reasoning and knowledge-based environment state descriptions. Integrating domain-specific knowledge into the prompt ensures semantically accurate and context-aware task plans. The proposed framework demonstrates its effectiveness by resolving semantic errors in symbolic plan generation, such as maintaining logical temporal goal ordering in scenarios involving hierarchical object placement. The proposed framework is validated through both simulation and real-world scenarios, demonstrating significant improvements over the baseline approach in terms of adaptability to dynamic environments and the generation of semantically correct task plans. Videos and code can be found here:  https://muhayyuddin.github.io/llm-tamp/.
@@ I Introduction
Task and motion planning (TAMP) is an essential component in sequential manipulation to perform tasks in complex environments. TAMP involves two stages: high-level symbolic plan generation, which develops a long-term abstract action sequence, and low-level motion planning, which determines the trajectories with geometric constraints required to execute the computed symbolic plan. The classical solution to TAMP problems consists of describing the task in a specific planning representation, such as the Planning Domain Description Language (PDDL) [1], and using a task planner, such as GraphPlan [2] or FastForward [3], to compute a symbolic plan, whose execution requires the call to a motion planner for all actions involving movements. A custom-designed communication layer is usually used to enable interaction between the task and motion planning modules.
Describing the task specification using PDDL for TAMP has proved to be very successful. However, PDDL has some drawbacks; it requires manual coding of all real-world descriptions and task constraints, which is impractical in a dynamic environment, and in addition, its syntax may be difficult for new users to understand [4]. On the other hand, compared to PDDL, natural language offers a more intuitive and user-friendly interface for task descriptions [5]. In this line, recent advances in pre-trained LLMs, such as GPT-4 [6], have shown good performance in various natural language tasks, and this has accelerated research on the use of LLMs in task planning [7] and TAMP [8] [9].
Figure 1: Task execution using the Onto-LLM-TAMP approach. The prompt is processed through the Prompt Tuning module, which uses knowledge-based reasoning to elaborate the user prompt. The state of the environment is textualized using ontological knowledge by the State Descriptor. This detailed elaboration enables the LLM to generate a semantically correct symbolic plan.
LLM-based methods for TAMP typically use pre-trained LLM models, such as GPT-4, that receive task descriptions as input [9]. The LLM module is responsible for computing symbolic plans and reasoning for failures, showing significant capabilities when in-context prompting is provided [10]. However, a weak prompt could drastically reduce planning and reasoning capabilities, and hence several recent studies proposed different ways of prompt engineering for better task elaboration. For example, a template-based prompt is proposed in [9], where fixed templates with variables are used to describe spatial and geometric relations, or the proposal in [8], where a template is used to describe the scene and another to textualize the state of the environment. To better exploit natural language, an interactive task planning using LLM was proposed in [11], where the user input was integrated with stored task guidelines to generate the prompt.
@@
To address the above-stated challenges, this study proposes an Onto-LLM-TAMP approach that utilizes knowledge-based reasoning to refine and expand the user’s prompt within the task context. Building on the baseline LLM3 TAMP framework [9], we extend it with our novel Onto-LLM-TAMP method. Some other approaches, such as [13], have also proposed to integrate ontological knowledge with LLM-based task planning, but their primary focus is on object properties to define optimal grasping strategies (e.g., glass $\rightarrow$ pick gently, wood $\rightarrow$ pick stably). On the contrary, we employ knowledge-based reasoning for prompt tuning to refine the action sequence for correct temporal goal ordering, integrating it with the user’s prompt input for the LLM. This ensures that the generated plan is not only correct but also semantically accurate. For example, consider the scenario depicted in Fig. 1, where the user instructs to ”put the banana, apple, and bowl in the plate”. In most cases, the symbolic baseline planner generates a sequence like ”pick banana and apple, and place in the plate” followed by ” pick bowl and place in the plate”. Although it is syntactically correct, it is semantically flawed, as it could result in placing the bowl on top of the apple and banana. Our proposed Onto-LLM-TAMP approach addresses this by incorporating domain-specific knowledge. It integrates a guiding note into the prompt, specifying that ”put bowl before food items because crockery has priority over food items. Put banana and apple after crockery because food items have less priority.” This extended prompt will help the LLM to generate a semantically correct and consistent plan.
Contributions
The core contribution of this work is a novel Onto-LLM-TAMP framework for task and motion planning that helps LLM to compute semantically accurate plans. Building upon this core framework, the following contributions are highlighted:
  • Automated prompt-tuning with task contextual reasoning: This approach extracts task-related common sense and procedural knowledge from the ontology to guide the LLM to generate a semantically accurate and logically ordered plan.
@@
  • Benchmarking of LLM models: We evaluated several major LLM models, including GPT, LLaMA, Gemini, and Cohere, in both standard LLM-TAMP and our Onto-LLM-TAMP frameworks to assess the impact of prompt tuning on the LLM performance.
The rest of the paper is structured as follows: Sec. II provides a detailed overview of the current state-of-the-art methods in Task and Motion Planning, highlighting the role of Large Language Models. Sec. III formulates the problem statement and presents the proposed Onto-LLM-TAMP framework. Sec. IV describes the proposed approach; elaborates on the methodology, and provides details of the ontology integration, and the execution pipeline. Sec. V presents experimental results and analysis, demonstrating the framework’s effectiveness in simulated and real-world environments. Finally, Sec. VI concludes the work.
@@ II-A Task and Motion Planning
TAMP uses a structured approach where task planning and motion planning are handled separately but in a coordinated manner. Task planning typically uses symbolic AI techniques, such as hierarchical planning, to generate a sequence of discrete actions based on predefined rules or logic. Motion planning uses geometric algorithms, such as sampling-based planners, to compute feasible paths [14] [15]. A declarative language, such as PDDL, is used for TAMP. However, its use is limited due to the requirement of a large number of parameters to define the problem, actions, and effects [16]. Moreover, as the action space expands, finding geometrically feasible symbolic action sequences becomes computationally challenging without effective heuristics [17]. Recent studies have explored data-driven heuristics to enhance TAMP efficiency [18] [19]. Other approaches use knowledge-based reasoning to reduce the action space, using domain-specific ontological knowledge, and then apply heuristics for efficient symbolic plan generation [20] [21]. The approach proposed here is a general manipulation framework to work across various domains, with the specific application ontology providing the specialized knowledge needed to correctly solve the task.
II-B Large Language Models
Large models represent a significant advancement in artificial intelligence, enabling powerful language understanding and general reasoning capabilities in various domains. Several models have been proposed including GPT-4 [22], Gemini [23], LLaMA [24], Cohere Language Models [25], and DeepSeek [26]. These models are built on transformer architectures trained on vast datasets. GPT-4 excels in complex reasoning and code synthesis, while Gemini integrates multimodal capabilities, allowing it to process text, images, and video. LLaMA stands out as an efficient open-source model that facilitates research and development by the broader community. Cohere offers industry-focused models that are optimized for industry-related applications. DeepSeek is another emerging model that aims to bridge the performance gap between open-source models and cutting-edge proprietary systems, focusing on multilingual and domain-specific tasks.
II-C LLMs for TAMP
@@
These foundation models are increasingly being used in task planning and symbolic reasoning for robotics and automated systems, i.e., LLMs are being explored to enhance task planning by using their ability to understand and generate human-like instructions, descriptions, and planning sequences. In particular, they allow to translate natural language task descriptions into formal task representations without the need to manually represent the problem domain [27, 28]. Due to their extensive language understanding and contextual reasoning capabilities, LLMs are used to generate high-level task plans and their translation into feasible motion sequences with minimal additional training [29, 30].
For example, GPT-4 has been adapted for robotic task planning by generating high-level action sequences and providing natural language explanations for task execution [31]. Google’s SayCan system combines LLMs with affordance-based planning to ground language commands into feasible robot actions [32]. Similarly, language models have been explored as zero-shot planners, enabling agents to extract actionable plans directly from textual instructions [33]. These approaches demonstrate the potential of foundation models to serve as cognitive reasoning engines in task and motion planning (TAMP). For instance, LLM-GROP [8] utilizes a pre-trained LLM to define symbolic goals for object placement and rearrangement, integrating these with traditional TAMP methods, and AutoTAMP [4] employs a pre-trained LLM to convert user inputs into formal language, which is then used in a TAMP algorithm.
@@
In all the approaches mentioned above, effective prompting techniques are crucial to optimize LLM performance [12]. In addition, [31] identified key challenges in prompting LLMs for robotic manipulation, including the need for detailed and accurate problem descriptions, as well as managing biases that shape response structures. This study tackles the issue of effective prompting for task and motion planning by utilizing ontological knowledge to enhance task elaboration for LLM.
Figure 2: The Onto-LLM-TAMP framework enhances prompt elaboration for generating semantically accurate symbolic plans. It begins by processing the user input to extract actions and objects through semantic tagging. The Contextual Inference Engine uses SPARQL queries to retrieve object types and priorities from the ontology, ensuring the correct action sequence based on predefined rules. The Perception Module, with YOLO-based object detection and FoundationPose for object pose estimation, provides real-time spatial data. This information is textualized using ontological knowledge by the Environmental State Descriptor and fed into the Prompt Generator. The final prompt is then fed into the LLM Task Planner, which produces a structured task plan. Finally, the Motion Planner ensures the robot executes the task with feasible, collision-free movements.
@@ III Problem Formulation
This section formally defines the task and motion planning problem and how we model the problem for Onto-LLM-TAMP.
@@ III-A Task and Motion Planning
  • $\mathcal{A}$ is the set of primitive actions, each one requiring the call to the motion planner for its execution. A primitive action $a\in\mathcal{A}$ is defined by a set of objects $o\in\mathcal{O}$ and a set of continuous parameters $\gamma$, e.g., Place(cup, [x, y, z, $\theta$]). The parameters $\gamma$ serve as goals for the motion planner and $a$ is considered feasible if the motion planner finds a solution, i.e., a collision-free trajectory $\tau$. We denote such a feasible action as $a(\tau)$.
@@
The objective in Task and Motion Planning is to derive a sequence of feasible actions $\mathbf{a}=\{a_{0}(\tau_{0}),a_{1}(\tau_{1}),\dots,a_{T}(\tau_{T})\}$, such that $s_{t+1}=\mathcal{T}(s_{t},a_{t}(\tau_{t}))$, for each $t=0,1,\dots,T$, and $g(s_{T+1})=1$.
@@ III-B Problem Modeling
The Onto-LLM-TAMP framework is defined as a tuple $\langle\mathcal{P},\mathcal{K},\xi,\mathcal{X},\mathcal{L}\rangle$, where each component plays a crucial role in structuring a feasible action plan:
@@
  • $\mathcal{K}$ denotes the ontological knowledge-base, which provides domain-specific information and task-related rules.
@@
    $\mathcal{L}:\mathcal{P}\times\epsilon\times\mathcal{X}\rightarrow\mathbf{a}$ (2)
This model receives the prompt $\mathcal{P}$, enhanced task knowledge $\epsilon$, and TAMP parameters $\mathcal{X}$, and it outputs a set of feasible actions $\mathbf{a}\subset\mathcal{A}$ that enable the robot to transition from the initial environment state to the desired goal state.
@@ IV Proposed approach
This section explains how ontological knowledge is used to generate suggestions for the LLM, ensuring semantically correct symbolic plan generation with proper temporal goal ordering. The architecture of our Onto-LLM-TAMP framework is illustrated in Fig. 2. It is composed of Ontological Prompt Construction Layer and the Planning Layer.
The Ontological Prompt Construction Layer is responsible for creating a prompt to be sent to the LLM-based planning module. This is achieved by using user input, predefined prompt templates, and the information available in an ontology. It comprises six modules: 1) the Ontology module that contains the application ontology, comprising classes, properties, and rules (detailed in Sec. IV-A); 2) the Semantic Tagging module that analyzes the user input and facilitates the queries to the ontology (discussed in Sec. IV-B); 3) the Contextual Inference Engine module that, using inferred knowledge composed of priorities and rules based on the relationships between objects and actions, suggests the right order of actions in the plan to perform the task in a semantically and logically correct manner (explained in Sec. IV-C); 4) the Perception Module that identifies all objects within the environment, estimating their spatial positions and orientations using YOLO [34] and FoundationPose [35], and instantiating them in the ontology, as done in a similar way in [36]; 5) the Environmental State Descriptor module that translates the environmental state retrieved from the ontology into a text format, describing the spatial layout and types of objects in the environment (the perception module and the environment state descriptor are detailed in Sec. IV-D); 6) The Prompt Generator module, that combines the prompt templates with the information coming from the Contextual Inference Engine and Environmental State Descriptor modules (explained in Sec. IV-E).
The Planning Layer takes a prompt as input and generates a plan to execute. It closely follows the baseline approach [9] and comprises three modules: 1): the User Input which describes the task to be performed; 2) the Prompt Template which contains the static sections of the prompt, provides general task descriptions and structures the overall prompt [^1] Complete template can be found here: https://github.com/Muhayyuddin/llm-tamp/blob/llm-site/assets/prompt/txt; 3) the LLM-TAMP Module, detailed in Sec. IV-F, that has the LLM Task Planner to compute a symbolic plan, and a motion planner to compute the motions for each action of the plan.
IV-A Ontological Knowledge Representation
@@
We developed a simple kitchen domain ontology to provide semantic reasoning for performing tasks in a kitchen environment. This domain-specific ontology enhances reasoning about task constraints, priority rules and object relations, and for this, three main classes have been defined: Task, ActionPriority and Object. The Object class is further subdivided into six subclasses describing different kitchen object types: FoodItems, BoxedFood, Crockery, Utensil, KitchenItems, and Container. The hierarchy of classes, object properties, data properties, and some example individuals are shown in Fig. 3.
@@
We define a set of rules [^2] Complete rule set used in this study is available here: https://github.com/Muhayyuddin/llm-tamp/blob/llm-site/assets/smart_kitchen.rdf that specify the correct order of action sequences when manipulating objects in the kitchen environment. These rules are implemented as instances of ActionPriority, where each rule associates a task (such as put or clean) with an object type (such as FoodItem or Crockery) and assigns a priority level. Through these rules, the reasoning mechanism infers a semantically correct action sequence for task execution. For example, Eq. (IV-A) and (IV-A) show the description logic representation for two rules, named Rule 1 and Rule 2.
@@
These rules define priority constraints for performing the ”Put” action on different object types. Rule 1 establishes that when placing objects, items classified as Crockery should have the highest priority, meaning they should be placed before other object types. This prioritization is explicitly set by assigning a priority value of 1 to Crockery, ensuring that these objects are handled first in the sequence of actions.
Similarly, Rule 2 specifies that objects that belong to the FoodItem category have a lower priority when performing the same ”Put” action. With a priority value of 2, FoodItems are placed after Crockery, reflecting a structured ordering in object manipulation. These rules collectively enforce a logical sequence for object handling, ensuring that tasks such as arranging or placing objects follow a predefined hierarchical structure that is suggested to LLM along with the task prompt.